OpenStudy (anonymous):

What is the sum of the arithmetic sequence 6, 14, 22 …, if there are 26 terms

OpenStudy (anonymous):

find the common difference first\[d = a_n - a_{n-1}\]and then use the sum formula for arithmetic progression\[S_n = \frac{n}{2}\left( 2a_{1} + \left( n - 1 \right)d \right)\]